# A Framework for Holistic KLD-based Waveform Design for Multi-User-Multi-Target ISAC Systems

### v4

@@ Abstract
This paper introduces a novel framework aimed at designing integrated waveforms for robust integrated sensing and communication (ISAC) systems. The system model consists of a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared-antenna deployment scenario. By leveraging Kullback-Leibler divergence (KLD) to holistically characterise both communication and sensing subsystems, three optimisation problems are formulated: (i) radar waveform KLD maximisation under communication constraints, (ii) communication waveform KLD maximisation subject to radar KLD requirements, and (iii) an integrated waveform KLD-based optimisation for ISAC that jointly balances both subsystems. The first two problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms. The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby strengthening the trade-off between sensing and communication objectives and achieving higher overall performance than the individual radar- or communication-only techniques. Unlike most existing ISAC waveform designs that regard communication signals solely as interference for sensing, the proposed framework utilises the holistic ISAC waveform—that is, the superimposed communication and sensing signals—to boost detection performance in the radar subsystem. Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming and identity-covariance radar baselines, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
@@ I Introduction
Integrated Sensing and Communications (ISAC) systems have gained significant interest as a key enabler of many 6G applications. ISAC aims to synergistically merge sensing and communication functionalities hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13]. This integration can be achieved through various approaches. For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilise resources. In contrast, the shared deployment strategy employs all antennas for both functionalities, hence offering higher resource efficiency, but at the cost of increased complexity due to the need for more sophisticated signal processing techniques[14, 15].
The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems, which significantly enhance both communication capacity and sensing accuracy [16]. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics—communication performance through achievable rate, outage probability, and bit error rate (BER), and sensing performance through estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimisation of ISAC systems. To this end, some unified performance measures have been introduced in the literature, such as the mutual information [18, 19, 20], and Kullback-Leibler divergence, also known as the relative entropy, which we recently introduced [21, 22, 23, 24].
@@ I-B Motivation and contribution
Despite these advances, current approaches often fall short of exploiting the full potential of the integrated system by assuming sensing and communication are competing subsystems rather than considering a synergic design. Most notably, there is a need for more comprehensive models that accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations, as shown in our paper, enhances radar detection capabilities, improving the system’s overall resource efficiency. This dual use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[31, 32].
Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC systems holistically, this paper proposes a novel ISAC waveform design with the objective of maximising KLD for the whole system as a design criterion in a shared antenna deployment scenario. Additionally, this work investigates the trade-off between trade-offs between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [22]. Our system model consists of a MIMO BS that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources, using a statistical radar model. To this end, the key contributions of this work are:
@@
  • Integrated Waveform Optimisation for ISAC: Beyond the individual radar and communication waveform-oriented designs, an integrated waveform optimisation framework for ISAC is introduced to optimise both waveforms together. This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, leading to enhanced performance in terms of radar detection accuracy and communication reliability compared with the separate waveform methods. By capturing and exploiting the synergy between sensing and data transmission, the unified technique achieves higher overall performance in shared-antenna ISAC deployments.
@@
  • Comprehensive Complexity Assessment: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility. This analysis is critical for determining practical implementation strategies in dynamic network environments where swift adaptation and efficient resource utilisation are essential.
Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline. The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, with modest radar performance gains. Both separate waveform techniques exhibit robust performance across varying SNR levels, with the radar waveform design demonstrating notably stable computational efficiency. The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering even higher overall performance in terms of sensing and communication signals. The numerical results highlight how leveraging communication signals for radar detection significantly enhances overall system performance, especially in scenarios with constrained dedicated radar resources. These findings emphasise the practical significance of KLD-based designs for shared-antenna ISAC systems, provide important adaptability and insights for balancing sensing and communication requirements in future wireless networks.
@@ I-C Paper organisation
The paper is structured as follows. Section II presents the system model. Sections III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system. In Section V, the radar waveform optimisation is investigated using the KLD. Section VI, the communication waveform optimisation is investigated using the KLD. Section VII shows the complexity analysis for both waveform optimisation techniques, Section VIII displays the numerical results, and finally, Section IX concludes the work.
$Notation$: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors. Superscripts ${(\cdot)}^{*}$, ${(\cdot)}^{T}$, and ${(\cdot)}^{H}$, denote the conjugate, transpose, and Hermitian transpose, respectively. Subscripts ${(\cdot)}_{\mathrm{c}}$ and ${(\cdot)}_{\mathrm{r}}$ relate to the communication, and radar subsystems, respectively.
@@ II System Model
We consider an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of ${T}$ targets and serving $K$ number of single-antenna communication UEs in the downlink direction in shared deployment manner. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ at the $l$-th snapshot can be formulated as follows,
$\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},$ (1)
@@
where $l\in\{1,2,\dots,L\}$ represents the discrete time index, and $L$ is the total number of snapshots considered for sensing outcome, $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets which reflect the received $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal. This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information. The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and .$\|.\|_{F}$ is the Frobenius norm.
@@ II-A Communication System
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. ZF designs which is typically used to precode the communication data symbols at the BS. This precoding scheme will serve as a benchmark for our designs. The ZF precoding eliminates inter-user interference (IUI) at the UEs regardless of the amount of the received power. This approach usually results in excellent performance at high SNR values. The normalised ZF beamforming matrix is designed also based on the channel matrix, and can be represented as follows,
where $\mathbf{\tilde{W}}_{\mathrm{c}}=\mathbf{H}^{H}\>\left(\mathbf{H}\>\mathbf{H}^{H}\right)^{-1}$is the ZF beamformer before normalisation, and $\mathbf{P}\in\mathbb{C}^{K\times K}=\mathrm{diag}\left(\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\right)$ is the power allocation matrix for the communication users, which is a diagonal matrix that controls the power allocated for each UE.

### v5

@@ Abstract
This paper introduces a novel framework aimed at designing integrated waveforms for robust integrated sensing and communication (ISAC) systems. The system model consists of a multiple-input multiple-output (MIMO) base station that simultaneously serves communication user equipments (UEs) and detects multiple targets using a shared-antenna deployment scenario. By leveraging Kullback-Leibler divergence (KLD) to holistically characterise both communication and sensing subsystems, three optimisation problems are formulated: (i) radar waveform KLD maximisation under communication constraints, (ii) communication waveform KLD maximisation subject to radar KLD requirements, and (iii) an integrated waveform KLD-based optimisation for ISAC that jointly balances both subsystems. The first two problems are solved using a projected gradient method with adaptive penalties for the radar waveforms and a gradient-assisted interior point method (IPM) for the communication waveforms. The third, integrated waveform optimisation approach adopts an alternating direction method of multipliers (ADMM) framework to unify radar and communication waveform designs into a single integrated optimisation, thereby synergising sensing and communication objectives and achieving higher overall performance than either radar- or communication-only techniques. Unlike most existing ISAC waveform designs that regard communication signals solely as interference for sensing, the proposed framework utilises the holistic ISAC waveform—that is, the superimposed communication and sensing signals—to boost detection performance in the radar subsystem. Simulation results show significant improvements in both radar detection and communication reliability compared with conventional zero-forcing beamforming, identity-covariance radar baselines, and traditional optimisation approaches, demonstrating the promise of KLD-based waveform designs for next-generation ISAC networks.
@@ I Introduction
Integrated Sensing and Communications (ISAC) systems have gained significant interest as a key enabler of many 6G applications. ISAC aims to synergistically merge sensing and communication functionalities, hence optimising base station (BS) resources for dual purposes [8, 9, 10, 11, 12, 13]. This integration can be achieved through various approaches. For example, in a separated deployment, the BS antennas are allocated distinctly between communication and radar subsystems, which can simplify the system design but may underutilise resources. In contrast, the shared deployment strategy employs all antennas for both functionalities, hence offering higher resource efficiency, but at the cost of increased complexity due to the need for more sophisticated signal processing techniques[14, 15].
The evolution of ISAC has been accelerated by advances in multi-antenna technologies, particularly massive Multiple-Input Multiple-Output (MIMO) systems, which significantly enhance both communication capacity and sensing accuracy [16]. However, evaluating the performance of ISAC systems poses unique challenges due to their dual functionality. Traditionally, communication and sensing subsystems have been assessed using distinct metrics—communication performance through achievable rate, outage probability, and bit error rate (BER), and sensing performance through estimation rate, detection probability, false alarm probability, and mean square error (MSE) [17]. This disparity complicates the holistic assessment and optimisation of ISAC systems. To this end, some unified performance measures have been introduced in the literature, such as mutual information [18, 19, 20, 21], and Kullback-Leibler divergence (KLD), also known as the relative entropy, which we recently introduced [22, 23, 24, 25]. Unlike mutual information approaches that assume capacity-achieving codes and focus on information transmission [26, 27], KLD directly measures distinguishability for practical finite-length systems, offers direct performance insights through Stein’s lemma for hypothesis testing, and captures the impact of practical modulation schemes on system performance, aligning naturally with ISAC’s dual objective.
@@ I-A Literature review
While the application of KLD in sensing systems is well-established [19, 28], its potential in characterising communication system performance has recently emerged as a powerful tool for ISAC design. For instance, in our previous works [23, 25], we employed KLD to analyse the performance trade-off in ISAC systems with separated antenna deployment, relating radar detection probability $P_{D}$ and communication BER to the achievable KLD. This connection stems from the fact that KLD is intrinsically tied to the maximum-likelihood detection, widely used in radar detection and equally central to communication receivers, whose errors are quantified by BER. Consequently, KLD can quantify the detection capability at both the radar and communication ends in a similar manner, making it a powerful and unified design tool for ISAC systems. While in [22] we proposed a low-complexity unified objective function based on KLD, the aim was focused on optimising network resources for both separated and shared deployment scenarios. Moreover, further insights about the achievable KLD trade-off in ISAC systems are explored in [29]. Nonetheless, the literature lacks a comprehensive investigation and beamforming design for a generalised multi-user-multi-target ISAC system using the promising KLD framework. The fundamental strength of KLD lies in its ability to provide a unified measure that captures the key detection processes in both subsystems. Unlike traditional metrics that operate in disparate mathematical spaces—probability of detection for radar and BER for communication, KLD enables direct comparison and joint optimisation. Moreover, KLD offers analytical tractability through closed-form gradient expressions, avoiding the computational complexity associated with Q-functions and error function derivatives in traditional formulations.
@@ I-B Motivation and contribution
Despite these advances, current approaches often fall short of exploiting the full potential of the integrated system by assuming sensing and communication are competing subsystems rather than considering a synergic design. Most notably, there is a need for more comprehensive models that accurately reflect the real-world operation of ISAC systems, particularly in scenarios where communication signals can be utilised for radar target detection alongside dedicated radar signals. Exploiting the interference from communication signals as a source of information for radar operations, as shown in our paper, enhances radar detection capabilities, improving the system’s overall resource efficiency. This dual use of communication signals represents a significant departure from traditional approaches, which typically treat interference as a detrimental factor to be mitigated rather than leveraged[36, 37]. Toward these objectives and motivated by the fact that KLD is able to provide a unified measure for ISAC systems holistically, this paper proposes a novel ISAC waveform design with the objective of maximising KLD for the whole system as a design criterion in a shared antenna deployment scenario. Additionally, this work investigates the synergies between sensing and communication in ISAC systems utilising KLD through well-established relations between KLD on one side and the detection probability in radar via Stein’s Lemma and BER in communication systems from the other side [23]. Our system model consists of a MIMO BS that simultaneously serves communication user equipment (UEs) and detects multiple targets using shared resources, using a statistical radar model. To this end, the key contributions of this work are:
@@
  • Integrated Waveform Optimisation for ISAC: Beyond the individual radar and communication waveform-oriented designs, an integrated waveform optimisation framework for ISAC is introduced to optimise both waveforms together. This integrated waveform approach leverages an alternating direction method of multipliers (ADMM) algorithm, yielding enhanced performance in radar detection accuracy and communication reliability compared with the separate waveform methods. By capturing and exploiting the synergy between sensing and data transmission, the unified approach achieves superior performance in shared-antenna ISAC deployments.
@@
  • Complexity Evaluation: We provide a thorough computational complexity analysis of our proposed ISAC algorithms, evaluating their scalability and real-time feasibility. This analysis is critical for determining practical implementation strategies in dynamic network environments where swift adaptation and efficient resource utilisation are essential.
  • Benchmark Analysis: We compare the proposed KLD-based strategies against the conventional zero-forcing (ZF) beamforming for the communication subsystem and the identity covariance design for the radar subsystem. In addition, we introduce two optimisation frameworks based on the BER and probability of detection to further benchmark our methods. These comparisons highlight the performance gains, computational complexity trade-offs, and practical relevance of our KLD-based approach in a range of ISAC scenarios.
Our results demonstrate that the proposed KLD-based optimisation techniques significantly outperform ZF beamforming and the radar subsystem’s identity covariance baseline, as well as the BER-based and detection-based designs in terms of computational cost. The radar waveform KLD-based optimisation achieves substantial gains in both target detection and communication quality, whereas the communication waveform KLD-based optimisation primarily benefits the communication subsystem, yielding modest gains in radar performance. Both separate waveform techniques exhibit robust performance across varying SNR levels, with the radar waveform design demonstrating notably stable computational efficiency. The integrated waveform optimisation KLD-based for ISAC approach unifies both radar and communication waveforms into a single integrated design, delivering higher overall performance in terms of sensing and communication signals. KLD-based formulations also facilitate a structured, scalable, and computationally efficient optimisation framework—unlike complex BER-based and detection-based designs, as evidenced by our numerical results. Notably, utilising the communication signals in radar detection significantly enhances overall system performance, especially in scenarios with limited dedicated radar resources. These highlight the practical significance of KLD-based waveform design and provide key insights for developing efficient shared deployment ISAC systems that balance sensing and communication requirements.
@@ I-C Paper organisation
The paper is structured as follows. Sec.II presents the system model. Sec.III and IV analyse the communication and radar systems, respectively, including the derivations of KLD for each system. In Sec.V, the radar waveform optimisation is investigated. Sec.VI, the communication waveform optimisation is investigated. Sec.VII, introduces the integrated waveform optimisation. Sec.VIII shows the complexity analysis for all waveform optimisation techniques, Sec.IX displays the numerical results, and finally, Sec.X concludes the work.
NotationNotation: Bold uppercase letters (e.g., $\mathbf{S}$) denote matrices, and bold lowercase letters (e.g., $\mathbf{s}$) denote vectors. Superscripts $(\cdot)$$^{*}$, ${(\cdot)}^{T}$, and ${(\cdot)}^{H}$, denote the conjugate, transpose, and Hermitian transpose, respectively. Subscripts ${(\cdot)}_{\mathrm{c}}$ and ${(\cdot)}_{\mathrm{r}}$ relate to the communication, and radar subsystems, respectively.
@@ II System Model
We consider an $N$ antenna MIMO-BS. The antennas are utilised for detecting a maximum number of T{T} targets and serving $K$ number of single-antenna communication UEs in the downlink direction in a shared deployment manner, reflecting the emerging paradigm where next-generation wireless infrastructure will inherently support dual-functional operation for both communication and sensing services [38]. The total transmitted power available at BS is $P_{\mathrm{T}}$ which is utilised for both sensing and data communication duties. The power $P_{\mathrm{T}}$ allocated to the radar and communication subsystems are denoted as $P_{\mathrm{r}}$ and $P_{\mathrm{c}}$, where $P_{\mathrm{T}}=P_{\mathrm{c}}+P_{\mathrm{r}}$. As a starting point, a non-optimised ZF beamforming technique is employed at BS to precode the information of communication UEs [16]. These techniques serve as benchmarks for our subsequent optimised systems, which will be introduced and analysed later in Sections V and VI. The combined transmitted ISAC signal $\mathbf{x}_{l}\in\mathbb{C}^{N\times 1}$ at the $l$-th snapshot can be formulated as follows,
$\mathbf{x}_{l}=\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}+\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r}},\vskip-3.61371pt$ (1)
@@
where $l\in\{1,2,\dots,L\}$ represents the discrete time index, and $L$ is the total number of snapshots considered for sensing outcome, $\mathbf{s}_{\mathrm{c},l}\in\mathbb{C}^{K\times 1}$ is a vector of communication UEs symbols, $\mathbf{s}_{\mathrm{r}}\in\mathbb{C}^{T\times 1}$ is a vector of the baseband radar waveforms for the potential targets, $\mathbf{W}_{\mathrm{c}}\in\mathbb{C}^{N\times K}$ represents the precoding matrix for the communication subsystem, and $\mathbf{W}_{\mathrm{r},l}\in\mathbb{C}^{N\times T}$ represents the precoding matrix for the radar subsystem. The transmitted signals vector $\mathbf{x}_{l}$ is transmitted from the BS and received at the UEs as well as the targets, which reflect the received $\mathbf{x}_{l}$ back to the BS. Although the radar signal component received at the UEs might cause interference that cannot be eliminated by the users, the communication signal component received at the targets and reflected back to the BS carries information about the target. Therefore, by carefully designing the precoding matrices at the BS, the sensing information can be enhanced by exploiting the communication signal. This dual-use approach enhances the radar performance by utilising the communication signal as an additional source of information. The radar waveform design is based on the combined transmit signal covariance matrix for the $t$-th beam where the covariance matrix of the $t$-th transmit beam is $\mathbf{R}_{t}=\frac{1}{L}\sum_{l=1}^{L}\mathbf{x}_{t,l}\>\mathbf{x}_{t,l}^{H}$, $\mathbf{x}_{t,l}=\mathbf{w}_{\mathrm{r},t,l}s_{\mathrm{r},t}+\mathbf{W}_{\mathrm{c}}\mathbf{s}_{\mathrm{c},l}$. This formulation allows us to incorporate the communication signals into the radar detection process. It should be noticed that the power for the radar is integrated into the precoding vector $\mathbf{w}_{\mathrm{r},t,l}$, where it is designed to have $\|\mathbf{w}_{\mathrm{r},t,l}\|_{F}^{2}=\frac{P_{\mathrm{r}}}{T}$, and $\mathbf{w}_{\mathrm{r},t,l}=\frac{P_{\mathrm{r}}}{NT}\>\mathbf{\tilde{w}}_{\mathrm{r},t,l}$, where $\mathbf{\tilde{w}}_{\mathrm{r},t,l}$ is the $t$th beam precoding vector, and $\|.\|_{F}$ is the Frobenius norm.
@@ II-A Communication System
where $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}\sim\mathcal{C}\mathcal{N}\left(0,\sigma_{h}^{2}\right)$ represents the channel from the BS to the $k$-th UE, $\omega_{\mathrm{MN},k}=\sum_{i=1,i\neq k}^{K}{\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{w}_{\mathrm{c},i}s_{\mathrm{c},i,l}}$ represents the IUI on the $k$-th UE from the other UEs, $d_{\mathrm{c},k}^{-\zeta/2}$ is the channel pathloss from BS to the $k$-th UE with $d_{\mathrm{c},k}$ representing the distance from BS to the $k$-th UE, and $\zeta$ is the pathloss exponent $\eta_{k}=\mathbf{h}_{k}^{T}\>d_{\mathrm{c},k}^{-\zeta/2}\>\mathbf{W}_{\mathrm{r},l}\mathbf{s}_{\mathrm{r},l}+n_{k,l}$ is the interference-plus-noise term with ${n}_{k,l}\sim\mathcal{C}\mathcal{N}\left(0,\>\sigma_{n}^{2}\right)$ representing the additive white Gaussian noise (AWGN). The precoding vector for the communication system is $\mathbf{w}_{\mathrm{c},k}\in\mathbb{C}^{N\times 1}$, which is typically designed based on the given channel matrix from MIMO-BS to UEs $\mathbf{H}\in\mathbb{C}^{N\times K}=\left[\mathbf{h}_{1},..,\mathbf{h}_{k},..,\mathbf{h}_{K}\right]$, where the elements of $\mathbf{H}$ are independent and identically distributed (i.i.d) complex Gaussian random variables with zero mean and variance $\sigma_{h}^{2}$. We assume that channel state information (CSI) is obtained through standard pilot-based estimation procedures widely adopted in modern communication systems with established techniques such as MMSE and ML estimators [39, 40].